# Modelling Cosmic-Ray Transport: Magnetized versus Unmagnetized Motion in Astrophysical Magnetic Turbulence

### v2

@@ 3.3 Model Results
Figure 10: Time-dependent MSDs for conditional and unconditional test-particle measurements, as well as conditional and combined model results. The combined model with optimized parameters shows good agreement with the unconditional test-particle MSD. The compound subdiffusion model is also shown for the unbiased optimized parameters, thus some deviation from the biased conditional data is present.
The MSDs of the combined model for the optimized values are plotted in figure 10 and agree well with the unconditional test-particle MSD for all considered particle energies. Note that the optimized magnetized mean durations are now comparable to the decoupling time scale, as shown in figure 9(a). Also plotted in figure 10 are the conditional MSDs and models, showing good agreement of the simple Langevin diffusion with the unmagnetized data, as well as some disagreement for the compound subdiffusion and the biased magnetized data. Specifically, as shown in figure 7(a), the optimization procedure predicts for most particle energies smaller MFPs compared to the conditional data. These smaller values can account for transport by less intense flux tubes or constellations of strongly correlated flux tubes, and thus correct the bias from the conditional magnetized data, dominated by isolated (intrinsic to the measurement methodology) and very intense flux tubes (fulfilling the heuristic magnetization criterion for long time scales).
@@ 4.1 Implications on Cosmic-Ray Transport
We emphasize that the considered regime of particle energies is non-asymptotic and limited by the available numerical resolution of the MHD simulation. However, in our results, consisting of MFPs in figure 7(a) and mean durations in figure 9(a), tendencies towards the asymptotic high- and low-energy regimes can be recognized. On one hand, the high-energy regime with $\hat{\omega}_{g}\lesssim 1$ will be entirely governed by unmagnetized motion, because particles with gyro radii $r_{g}\gtrsim L_{u}$ average out any non-trivial structures, as shown by increasing $t^{\ast}_{\mathrm{scatter}}/t^{\ast}_{\mu}$ and $\lambda_{\mathrm{asymp}}/\lambda_{\mathrm{scatter}}\to 1$ as $\hat{\omega}_{g}$ decreases. The expected random-walk scaling is $\lambda\sim\hat{\omega}_{g}^{-2}$ (Reichherzer et al. 2020; the indicated scaling $\lambda\sim\hat{\omega}_{g}^{-1}$ in figure 7(a) is only transitional, compare also Lübke et al. 2024). The conditional magnetized model is valid to roughly $\hat{\omega}_{g}\sim 64$.
On the other hand, the low-energy regime $\hat{\omega}_{g}\gg 1$ becomes more and more dominated by magnetized motion as indicated by decreasing $t^{\ast}_{\mathrm{scatter}}/t^{\ast}_{\mu}$. The asymptotic MFP $\lambda_{\mathrm{asymp}}$ likely becomes a function of the coherent field line MFP $\lambda_{\mathrm{fl}}$ and the large-scale mirror MFP $\lambda_{\mu}$ (our previous assumption $\lambda_{\mathrm{fl}}=\lambda_{\mu}$ is not expected to hold in general), as well as the mean magnetized duration $t^{\ast}_{\mu}$, which indicates the typical particle time scale of decoupling from field lines. The scaling $\lambda\sim\hat{\omega}_{g}^{-1/3}$ indicated in figure 7(a) serves purely as reference and should be taken with care, as small gyro radii with $\hat{\omega}_{g}\gtrsim 256$ are significantly affected by the MHD dissipation scales (compare also figure 1).
Our results clearly highlight the crucial role of coherent magnetic structures, especially for the low-energy regime, and support the role of field-line wandering in cosmic ray transport, as discussed by Pezzi & Blasi (2024). Further, highly-magnetized low-energy particles such as GeV galactic cosmic rays are expected to be dynamically relevant on the plasma dynamics, requiring the consideration of streaming (Sampson et al., 2023).
In order to speculate about the asymptotic behaviour of the parameters of magnetized motion, we presume that the field line MFP $\lambda_{\mathrm{fl}}$ for a particle with $\hat{\omega}_{g}$ is given by the average coherent curvature radius, i.e., the average of curvature radii larger than the typical gyro radius $\frac{\pi}{2}\hat{\omega}_{g}^{-1}$,
@@
which agrees well with our MHD simulation for $s=2$ and $\kappa_{0}=1.5$, and yields $\lim_{\hat{\omega}_{g}\to\infty}\lambda_{\mathrm{fl}}(\hat{\omega}_{g})=\kappa_{0}^{-1}$.
Further, the path lengths between large-scale mirror events likely follow a broad distribution, which is truncated by the typical length of coherent field lines, so we remain with our assumption $\lambda_{\mu}\approx\lambda_{\mathrm{fl}}$ for now. Given constant field-line and mirroring MFPs $\lambda_{\mathrm{fl}}\approx\lambda_{\mu}\approx\kappa_{0}^{-1}$ in the low-energy limit, the asymptotic MFP is then solely determined by the decoupling time scale $t^{\ast}_{\mu}$, which in combination with compound subdiffusion of particles along coherent field lines predicts $\lambda_{\mathrm{asymp}}\sim{t^{\ast}_{\mu}}^{-1/2}$. As shown in figure 8(a), the durations of magnetized motion follow a broad power-law distribution with a long-time cut-off, which increases with decreasing particle energy. The asymptotic scaling of $t^{\ast}_{\mu}$ is not resolved in our data and can only be speculated about at the moment. For instance, $\lambda_{\mathrm{asymp}}\sim\hat{\omega}_{g}^{-1/3}$ would suggest $t^{\ast}_{\mu}\sim\hat{\omega}_{g}^{2/3}$. We expect it to arise from a complex interplay of lengths of coherent structures $\lambda_{\mathrm{fl}}$, mirroring rates $c/\lambda_{\mu}$, drift motions, and possibly gyro-resonance due to Alfvén waves travelling along coherent flux tubes.
@@ 4.2 Discussion of Coherent Structures
Despite the idealized nature of our turbulence setup, consisting of a visco-resistive incompressible MHD simulation with $Re_{T}\sim O(100)$ and $Pr={\nu}/{\eta}=1$, our results can be contextualized with coherent structures ubiquitously found in simulations and observations (see, e.g., Robitaille et al., 2020; Ntormousi et al., 2024). The fluctuating dynamo with compressible (Beattie et al., 2024) or kinetic (St-Onge & Kunz, 2018) physics also exhibits pronounced coherent structures, although their statistics, such as $p(B,\kappa)$, likely differ from our case and a dedicated study of cosmic-ray transport in these cases would be useful. While our focus has been on intense flux tubes generated by the fluctuating dynamo, coherent structures also emerge as patch-wise aligned states due to selective decay (Matthaeus et al. 2015; Hosking & Schekochihin 2021; see also figure 3) or as intense aligned wave packets in critically-balanced turbulence (Perez & Boldyrev, 2009; Chandran et al., 2015). Generally, different turbulence setups (e.g., forced or freely decaying, with or without an imposed background magnetic field, different degrees of compressibility, cross- and magnetic helicity) lead to different predominant kinds of structures, which may impact cosmic-ray transport in different ways. Here we use the term “structure” to loosely summarize relatively coherent phenomena with non-linear qualities (see also Grošelj et al. 2019 for a different but not contradictory point of view).
@@
Current sheets emerge as the unifying feature of these various kinds of turbulence (see, e.g., Servidio et al., 2011; Zhdankin et al., 2016), which, given sufficient dynamical range, become tearing-unstable, leading to fast reconnection and the seeding of small-scale plasmoid chains, i.e., complexes of highly-tangled small-scale flux tubes. This process is associated with a reconnection-mediated turbulence regime at the transition to kinetic scales (Loureiro & Boldyrev, 2017; Mallet et al., 2017; Chernoglazov et al., 2021; Dong et al., 2022), which has important implications on cosmic-ray acceleration (Comisso & Sironi, 2019) and may be relevant for a transport theory of the lowest-energy cosmic rays or suprathermal particle populations. However, particular care must be taken when analysing simulations featuring plasmoids, such as ours, because they may also appear as numerical artifacts for insufficient numerical resolution (Morillo & Alexakis, 2025).
@@
Understanding these various structures and their interaction with charged particles is crucial for a comprehensive picture of cosmic-ray transport in realistic multiphase media (Armillotta et al., 2022; Beattie et al., 2025). Also, given sufficiently high numerical resolution, gyro-resonance can be expected to re-appear, for instance mediated by small-scale Alfvén wave turbulence emerging along coherent flux tubes, either as part of the turbulent cascade or self-generated by cosmic rays. A new cosmic-ray transport theory should carefully weigh these different processes and their varying predominance against each other (Kempski & Quataert, 2022; Hopkins et al., 2022).
@@ 4.3 Synthetic Turbulence
Synthetic turbulence refers to a class of algorithms designed for overcoming the limited resolvable range of scales of first-principles simulations. They provide fast generation of random fields resembling certain statistics expected for realistic turbulence. Most commonly, guided by the paradigm of gyro-resonance, these fields are synthesized from a prescribed energy spectrum and random Fourier phases (see, e.g., Mertsch, 2020). These random-phase models have well-known shortcomings, such as the inability to reproduce intermittency and coherent structures, as illustrated by, e.g., Shukurov et al. (2017), which has led to the exploration of intermittent and structured models (Subedi et al., 2014; Pucci et al., 2016; Durrive et al., 2022; Lübke et al., 2024; Maci et al., 2024; Lesaffre et al., 2025). To address the crucial geometric structure, future synthetic models for strong isotropic turbulence should include the field-line curvature distribution $p(\kappa)\sim\kappa^{-5/2}$ as well as the anti-correlation between magnetic field strength and field-line curvature $B\sim\kappa^{-1/2}$. Instead of generating a synthetic field directly, Generative Diffusion Models, a recently developed machine learning technique, are able to learn and synthesize particle trajectories directly (Li et al., 2023; Martin et al., 2025). Such a technique could also be useful for classifying the various transport behaviours and refining the heuristic magnetization criterion $\bar{\kappa}\bar{r}_{g}\sim 1$.
@@ 5.1 Refinement of the Combined Stochastic Model
We emphasize that our model is semi-empirical, where parameters are inferred from the test-particle data. It is essential for a proper transport theory to predict the involved MFPs and mean durations from the underlying turbulence. In particular, such a theory should account for the hierarchy of scales present in the problem, from scattering on gyro scales, over motion within individual coherent structures, and up to large-scale network-like constellations of the population of coherent structures (Ntormousi et al., 2024). This task could be approached by projector-based coarse-graining of the dynamics, following the Mori-Zwanzig formalism (Hudson & Li, 2020; Lin et al., 2021).
Another possible approach to accurately derive the required MFPs and mean durations would be to follow Drummond (1982) and start with a path integral formulation of turbulent diffusion. However, in Drummond’s approach the magnetic potential is assumed to be Gaussian. A possible, far from trivial approach would be a path integral formulation including the action $S_{\mathrm{MHD}}$ corresponding to the MHD dynamics plus the action $S_{\mathrm{particle}}$ of the particle dynamics and applying non-perturbative methods to approximate this complex path integral (see, e.g., Grafke et al., 2015; Bureković et al., 2024; Schorlepp & Grafke, 2025).
@@ 5.2 Alternative Transport Descriptions
A central assumption of our model is that all involved kinds of motion are Gaussian and sufficiently characterized by their MFPs or MSDs. While our combined model of Langevin and compound subdiffusion successfully reproduces the test-particle MSD, possibly relevant higher-order statistics are neglected. For instance, intermittent pitch-angle scattering (Zimbardo & Perri, 2020) or streaming (Sampson et al., 2023) can lead to superdiffusion of cosmic rays. In the context of our test-particle data, magnetized motion may be described by spatial superdiffusion (averaged over the pitch angle), and mirror-confinement in coherent structures may correspond to subdiffusion. A clever combination of Lévy walks and extended waiting times (Zaburdaev et al., 2015) with truncated distributions (Liang & Oh, 2025) may reproduce the intricate time evolution of the test-particle MSDs, while also accounting for possibly relevant higher-order statistics. In this case of competition between sub- and superdiffusion (Magdziarz & Weron, 2007) of cosmic rays, a single (fractional) transport equation might be formulated that is equivalent to the stochastic model as presented for the superdiffusive case in Effenberger et al. (2024) and Aerdker et al. (2025).
@@
Alternatively, especially when modelling large-scale ($\gg L_{u}$) multiphase media, the structured nature of turbulence may be reflected by distinct diffusion coefficients. For the example of two spatially distinguishable phases, one can employ a regular Fokker-Planck equation $\partial_{t}f(\boldsymbol{x},t)=\Delta D(\boldsymbol{x})f(\boldsymbol{x},t)$ with a mixed diffusion coefficient $D(\boldsymbol{x})=D_{1}$ for $\boldsymbol{x}\in\text{phase 1}$ and $D(\boldsymbol{x})=D_{2}$ for $\boldsymbol{x}\in\text{phase 2}$, which is readily applicable to astrophysical problems (see, e.g., Reichherzer et al., 2025). Otherwise, a temporal switching Fokker-Planck equation $\partial_{t}f_{i}(\boldsymbol{x},t)=D_{i}\Delta f_{i}(\boldsymbol{x},t)+A_{ij}f_{j}(\boldsymbol{x},t)$ with Markovian switching rates $A_{ij}$ may be useful (see, e.g., Bressloff & Lawley, 2017; Balcerek et al., 2023), which could be applied as a sub-grid model for cosmic-ray transport.
@@
All presented models involve a competition between two transport regimes to capture the transient phases of the observed MSD. Similar switching or competing models are also known in other complex systems (see, e.g., Doerries et al., 2022; Datta et al., 2024).
5.3 Generalized Transport Equation

### v3

@@ 3.3 Model results
Figure 10: Time-dependent MSDs for conditional and unconditional test-particle measurements, as well as conditional and combined model results. The combined model with optimised parameters shows good agreement with the unconditional test-particle MSD. The compound subdiffusion model is also shown for the unbiased optimised parameters, thus some deviation from the biased conditional data is present.
The MSDs of the combined model for the optimised values are plotted in figure 10 and agree well with the unconditional test-particle MSD for all considered particle energies. Note that the optimised magnetised mean durations are now comparable to the decoupling time scale, as shown in figure 9(a). Also plotted in figure 10 are the conditional MSDs and models, showing good agreement of the simple Langevin diffusion with the unmagnetised data, as well as some disagreement for the compound subdiffusion and the biased magnetised data. Specifically, as shown in figure 7(a), the optimisation procedure predicts for most particle energies smaller MFPs compared to the conditional data. These smaller values can account for transport by less intense flux tubes or constellations of strongly correlated flux tubes and, thus, correct the bias from the conditional magnetised data, dominated by isolated (intrinsic to the measurement methodology) and very intense flux tubes (fulfilling the heuristic magnetisation criterion for long time scales).
@@ 4.1 Implications on cosmic-ray transport
We emphasise that the considered regime of particle energies is non-asymptotic and limited by the available numerical resolution of the MHD simulation. However, in our results, consisting of MFPs in figure 7(a) and mean durations in figure 9(a), tendencies towards the asymptotic high- and low-energy regimes can be recognised. On one hand, the high-energy regime with $\hat{\omega}_{g}\lesssim 1$ will be entirely governed by unmagnetised motion, because particles with gyro radii $r_{g}\gtrsim L_{u}$ average out any non-trivial structures, as shown by increasing $t^{\ast}_{\mathrm{scatter}}/t^{\ast}_{\mu}$ and $\lambda_{\mathrm{asymp}}/\lambda_{\mathrm{scatter}}\to 1$ as $\hat{\omega}_{g}$ decreases. The expected random-walk scaling is $\lambda\sim\hat{\omega}_{g}^{-2}$ (Reichherzer et al. 2020; the indicated scaling $\lambda\sim\hat{\omega}_{g}^{-1}$ in figure 7(a) is only transitional, compare also Lübke et al. 2024). The conditional magnetised model is valid to roughly $\hat{\omega}_{g}\sim 64$.
On the other hand, the low-energy regime $\hat{\omega}_{g}\gg 1$ becomes increasingly dominated by magnetised motion as indicated by decreasing $t^{\ast}_{\mathrm{scatter}}/t^{\ast}_{\mu}$. The asymptotic MFP $\lambda_{\mathrm{asymp}}$ likely becomes a function of the coherent field line MFP $\lambda_{\mathrm{fl}}$ and the large-scale mirror MFP $\lambda_{\mu}$ (our previous assumption $\lambda_{\mathrm{fl}}=\lambda_{\mu}$ is not expected to hold in general), as well as the mean magnetised duration $t^{\ast}_{\mu}$, which indicates the typical particle time scale of decoupling from field lines. The scaling $\lambda\sim\hat{\omega}_{g}^{-1/3}$ indicated in figure 7(a) serves purely as reference and should be taken with care, as small gyro radii with $\hat{\omega}_{g}\gtrsim 256$ are significantly affected by the MHD dissipation scales (compare also figure 1).
Our results clearly highlight the crucial role of coherent magnetic structures, especially for the low-energy regime, and support the role of field-line wandering in cosmic ray transport, as discussed by Pezzi & Blasi (2024). Further, highly magnetised low-energy particles such as GeV galactic cosmic rays are expected to be dynamically relevant on the plasma dynamics, requiring the consideration of streaming (Sampson et al., 2023).
To speculate about the asymptotic behaviour of the parameters of magnetised motion, we presume that the field line MFP $\lambda_{\mathrm{fl}}$ for a particle with $\hat{\omega}_{g}$ is given by the average coherent curvature radius, i.e. the average of curvature radii larger than the typical gyro radius $\frac{\pi}{2}\hat{\omega}_{g}^{-1}$,
@@
which agrees well with our MHD simulation for $s=2$ and $\kappa_{0}=1.5$ and yields $\lim_{\hat{\omega}_{g}\to\infty}\lambda_{\mathrm{fl}}(\hat{\omega}_{g})=\kappa_{0}^{-1}$.
Further, the path lengths between large-scale mirror events likely follow a broad distribution, which is truncated by the typical length of coherent field lines, so we remain with our assumption $\lambda_{\mu}\approx\lambda_{\mathrm{fl}}$ for now. Given constant field-line and mirroring MFPs $\lambda_{\mathrm{fl}}\approx\lambda_{\mu}\approx\kappa_{0}^{-1}$ in the low-energy limit, the asymptotic MFP is then solely determined by the decoupling time scale $t^{\ast}_{\mu}$, which, in combination with compound subdiffusion of particles along coherent field lines, predicts $\lambda_{\mathrm{asymp}}\sim{t^{\ast}_{\mu}}^{-1/2}$. As shown in figure 8(a), the durations of magnetised motion follow a broad power-law distribution with a long-time cut-off, which increases with decreasing particle energy. The asymptotic scaling of $t^{\ast}_{\mu}$ is not resolved in our data and can only be speculated about at the moment. For instance, $\lambda_{\mathrm{asymp}}\sim\hat{\omega}_{g}^{-1/3}$ would suggest $t^{\ast}_{\mu}\sim\hat{\omega}_{g}^{2/3}$. We expect it to arise from a complex interplay of lengths of coherent structures $\lambda_{\mathrm{fl}}$, mirroring rates $c/\lambda_{\mu}$, drift motions and possibly gyro-resonance due to Alfvén waves travelling along coherent flux tubes.
@@ 4.2 Discussion of coherent structures
Despite the idealised nature of our turbulence set-up, consisting of a visco-resistive incompressible MHD simulation with $Re_{T}\sim O(100)$ and $Pr={\nu}/{\eta}=1$, our results can be contextualised with coherent structures ubiquitously found in simulations and observations (see, e.g. Robitaille et al., 2020; Ntormousi et al., 2024). The fluctuating dynamo with compressible (Beattie et al., 2024) or kinetic (St-Onge & Kunz, 2018) physics also exhibits pronounced coherent structures, although their statistics, such as $p(B,\kappa)$, likely differ from our case and a dedicated study of cosmic-ray transport in these cases would be useful. While our focus has been on intense flux tubes generated by the fluctuating dynamo, coherent structures also emerge as patch-wise aligned states due to selective decay (Matthaeus et al. 2015; Hosking & Schekochihin 2021; see also figure 3) or as intense aligned wave packets in critically balanced turbulence (Perez & Boldyrev, 2009; Chandran et al., 2015). Generally, different turbulence set-ups (e.g. forced or freely decaying, with or without an imposed background magnetic field, different degrees of compressibility, cross- and magnetic helicity) lead to different predominant kinds of structures, which may impact cosmic-ray transport in different ways. Here, we use the term “structure” to loosely summarise relatively coherent phenomena with nonlinear qualities (see also Grošelj et al. 2019 for a different, but not contradictory, point of view).
@@
Current sheets emerge as the unifying feature of these various kinds of turbulence (see, e.g. Servidio et al., 2011; Zhdankin et al., 2016), which, given sufficient dynamical range, become tearing-unstable, leading to fast reconnection and the seeding of small-scale plasmoid chains, i.e. complexes of highly-tangled small-scale flux tubes. This process is associated with a reconnection-mediated turbulence regime at the transition to kinetic scales (Loureiro & Boldyrev, 2017; Mallet et al., 2017; Chernoglazov et al., 2021; Dong et al., 2022), which has important implications on cosmic-ray acceleration (Comisso & Sironi, 2019) and may be relevant for a transport theory of the lowest-energy cosmic rays or suprathermal particle populations. However, particular care must be taken when analysing simulations featuring plasmoids, such as ours, because they may also appear as numerical artefacts for insufficient numerical resolution (Morillo & Alexakis, 2025).
@@
Understanding these various structures and their interaction with charged particles is crucial for a comprehensive picture of cosmic-ray transport in realistic multiphase media (Armillotta et al., 2022; Beattie et al., 2025). Also, given sufficiently high numerical resolution, gyro-resonance can be expected to re-appear, for instance, mediated by small-scale Alfvén wave turbulence emerging along coherent flux tubes, either as part of the turbulent cascade or self-generated by cosmic rays. A new cosmic-ray transport theory should carefully weigh these different processes and their varying predominance against each other (Kempski & Quataert, 2022; Hopkins et al., 2022).
@@ 4.3 Synthetic turbulence
Synthetic turbulence refers to a class of algorithms designed for overcoming the limited resolvable range of scales of first-principles simulations. They provide fast generation of random fields resembling certain statistics expected for realistic turbulence. Most commonly, guided by the paradigm of gyro-resonance, these fields are synthesised from a prescribed energy spectrum and random Fourier phases (see, e.g. Mertsch, 2020). These random-phase models have well-known shortcomings, such as the inability to reproduce intermittency and coherent structures, as illustrated by, e.g. Shukurov et al. (2017), which has led to the exploration of intermittent and structured models (Subedi et al., 2014; Pucci et al., 2016; Durrive et al., 2022; Lübke et al., 2024; Maci et al., 2024; Lesaffre et al., 2025). To address the crucial geometric structure, future synthetic models for strong isotropic turbulence should include the field-line curvature distribution $p(\kappa)\sim\kappa^{-5/2}$ as well as the anti-correlation between magnetic field strength and field-line curvature $B\sim\kappa^{-1/2}$. Instead of generating a synthetic field directly, Generative Diffusion Models, a recently developed machine learning technique, are able to learn and synthesise particle trajectories directly (Li et al., 2023; Martin et al., 2025). Such a technique could also be useful for classifying the various transport behaviours and refining the heuristic magnetisation criterion $\bar{\kappa}\bar{r}_{g}\sim 1$.
@@ 5.1 Refinement of the combined stochastic model
We emphasise that our model is semi-empirical, where parameters are inferred from the test-particle data. It is essential for a proper transport theory to predict the involved MFPs and mean durations from the underlying turbulence. In particular, such a theory should account for the hierarchy of scales present in the problem, from scattering on gyro scales, over motion within individual coherent structures and up to large-scale network-like constellations of the population of coherent structures (Ntormousi et al., 2024). This task could be approached by projector-based coarse-graining of the dynamics, following the Mori-Zwanzig formalism (Hudson & Li, 2020; Lin et al., 2021).
Another possible approach to accurately derive the required MFPs and mean durations would be to follow Drummond (1982) and start with a path integral formulation of turbulent diffusion. However, in Drummond’s approach, the magnetic potential is assumed to be Gaussian. A possible, far from trivial approach would be a path integral formulation including the action $S_{\mathrm{MHD}}$ corresponding to the MHD dynamics plus the action $S_{\mathrm{particle}}$ of the particle dynamics and applying non-perturbative methods to approximate this complex path integral (see, e.g. Grafke et al., 2015; Bureković et al., 2024; Schorlepp & Grafke, 2025).
@@ 5.2 Alternative transport descriptions
A central assumption of our model is that all involved kinds of motion are Gaussian and sufficiently characterised by their MFPs or MSDs. While our combined model of Langevin and compound subdiffusion successfully reproduces the test-particle MSD, possibly relevant higher-order statistics are neglected. For instance, intermittent pitch-angle scattering (Zimbardo & Perri, 2020) or streaming (Sampson et al., 2023) can lead to superdiffusion of cosmic rays. In the context of our test-particle data, magnetised motion may be described by spatial superdiffusion (averaged over the pitch angle), and mirror-confinement in coherent structures may correspond to subdiffusion. A clever combination of Lévy walks and extended waiting times (Zaburdaev et al., 2015) with truncated distributions (Liang & Oh, 2025) may reproduce the intricate time evolution of the test-particle MSDs, while also accounting for possibly relevant higher-order statistics. In this case of competition between sub- and superdiffusion (Magdziarz & Weron, 2007) of cosmic rays, a single (fractional) transport equation might be formulated that is equivalent to the stochastic model as presented for the superdiffusive case in Effenberger et al. (2024) and Aerdker et al. (2025).
@@
Alternatively, especially when modelling large-scale ($\gg L_{u}$) multiphase media, the structured nature of turbulence may be reflected by distinct diffusion coefficients. For the example of two spatially distinguishable phases, one can employ a regular Fokker-Planck equation $\partial_{t}f(\boldsymbol{x},t)=\Delta D(\boldsymbol{x})f(\boldsymbol{x},t)$ with a mixed diffusion coefficient $D(\boldsymbol{x})=D_{1}$ for $\boldsymbol{x}\in\text{phase 1}$ and $D(\boldsymbol{x})=D_{2}$ for $\boldsymbol{x}\in\text{phase 2}$, which is readily applicable to astrophysical problems (see, e.g. Reichherzer et al., 2025). Otherwise, a temporal switching Fokker-Planck equation $\partial_{t}f_{i}(\boldsymbol{x},t)=D_{i}\Delta f_{i}(\boldsymbol{x},t)+A_{ij}f_{j}(\boldsymbol{x},t)$ with Markovian switching rates $A_{ij}$ may be useful (see, e.g. Bressloff & Lawley, 2017; Balcerek et al., 2023), which could be applied as a sub-grid model for cosmic-ray transport.
@@
All presented models involve a competition between two transport regimes to capture the transient phases of the observed MSD. Similar switching or competing models are also known in other complex systems (see, e.g. Doerries et al., 2022; Datta et al., 2024).
5.3 Generalised transport equation